OpenStudy (anonymous):

Question- Find the common difference of each arithmetic sequence? Numbers given: -3, -0.6, 1.8, 4.2?

OpenStudy (anonymous):

Difference means the outcome of subtraction, as you know. Subtract the sequential terms. -3 - -0.6 = -3 + 0.6 = -(3 - 0.6) = -(2.4) = -2.4 -0.6 - 1.8 = -1.8 - 0.6 = -(1.8 + 0.6) = -(2.4) = -2.4 1.8 - 4.2 = -4.2 + 1.8 = -(4.2 - 1.8) = -(2.4) = -2.4

OpenStudy (anonymous):

Thanks so much, so helpful!!

OpenStudy (anonymous):

Since it's given to you that it's an arithmetic sequence, you could actually just do one of those. However, if it's not given, you would have to do all 3 to verify that it's an arithmetic sequence. And you're welcome.